# Debiased machine learning for combining probability and non-probability survey data

### v1

@@ 1 Introduction
Several methods have recently been proposed for using data from a probability sample (Sample A) as auxiliary information to address nonrepresentative of a nonprobability sample (Sample B). These methods fall into three classes, all of which require data on some covariates $X$ that are observed on the individuals in both samples. The first class of methods, which include inverse probability weighting (IPW) methods, involve estimating weights for the individuals in Sample B such that the weighted Sample B is representative of the population. The second class, known as mass imputation methods, involve using the relation between $X$ and $Y$ in Sample B to impute the $Y$ values in Sample A. The third class, known as double robust (DR) methods, combine IPW and mass imputation. IPW methods require a consistent estimator of the selection probability function for Sample B, i.e. the conditional probability that an individual belongs to Sample B given $X$. Mass imputation methods require a consistent estimator of the conditional expectation of $Y$ given $X$. We shall refer to this conditional probability and expectation as ‘nuisance’ functions, because they are not of direct interest. DR methods use estimators of both nuisance functions but only require one of these estimators to be consistent.
@@
A parametric model could be used for each of the two nuisance functions, but the resulting estimator would typically be inconsistent if that parametric model — or, in the case of DR estimators, both parametric models — were misspecified. There is therefore considerable interest in using more flexible ‘data-adaptive’ or ‘machine-learning’ estimators of the nuisance functions, in order to minimise this risk of model misspecification. A number of researchers have done this. Ferri-Garcia et al. (2020, 2024)[16, 17], Rueda et al. (2023)[13] and Rueda et al. (2024)[12] use IPW and estimate the conditional probability nuisance function with k-nearest neighbours, random forest or XGboost. Castro-Martin et al. (2020)[2] use IPW or mass imputation, and estimate the conditional probability or expectation nuisance function using random forest or XGboost. Ferri-Garcia et al. (2022)[15] use IPW and estimate the nuisance function using classification and regression trees (CART). Castro-Martin et al. (2021)[3] use IPW, mass imputation or a DR estimator, and estimate the nuisance functions (or functions) using XGboost. Cobo et al. (2025)[11] use a DR estimator and estimate the nuisance functions using XGboost, k-nearest neighbours or neural nets. To reduce the variability of the estimated inverse probability weights, Castro-Martin et al. (2021)[3], Rueda et al. (2024)[12] and Cobo et al. (2025)[11] apply kernel smoothing to them. Wang et al. (2020,2022)[27, 28] developed a theoretical justification for this kernel smoothing approach when logistic regression is used, but not when data-adaptive methods are used. Chen et al. (2022)[5] use mass imputation with kernel smoothing, show that the resulting estimator of $\bar{Y}$ is asymptotically normal and provide a formula for the variance. They note, however, that this method is subject to the curse of dimensionality.
@@
A closely related problem to that of estimating the mean of $Y$ in a finite population by combining data from probability and nonprobability samples is that of estimating $E(Y)$, the expectation of $Y$ in an infinite population, when data $(X_{i},Y_{i})$ $(i=1,\ldots,n)$ are independent and identically distributed (iid), $X$ is fully observed and $Y$ is missing at random given $X$. IPW, imputation and DR methods exist for this problem. The first two require consistent estimators of, respectively, the conditional probability that $Y$ is observed given $X$ and the conditional expectation of $Y$ given $X$. DR methods involve estimating both these nuisance functions, but require only that (at least) one of them is estimated consistently. For IPW and imputation, it is known that using data-adaptive methods to estimate the nuisance function can cause problems. Data-adaptive methods typically yield nuisance function estimators that converge slowly, and this slow convergence is inherited by the IPW or imputation estimator of $E(Y)$, which can then have considerable finite-sample bias; confidence intervals can also have poor coverage[26]. The DR approach (also known as ‘debiased machine learning’) can overcome this issue, because DR estimators do not inherit the slow convergence of the nuisance function estimators[26, 14, 23, 20, 8, 7, 9]. Given the similarity of the problem of estimating $E(Y)$ for an infinite population with iid data to the problem of estimating the mean of $Y$ in a finite population using a nonprobability sample, there is reason to be concerned that the same issue may well also apply to the latter problem. In this article, we study the use of a DR estimator of the mean of $Y$ in a finite population when combining data from a probability sample and a nonprobability sample and using data-adaptive estimators of the nuisance functions. Our goal is to provide a theoretical justification for the validity of this approach. An important difference between the problem of estimating $E(Y)$ for an infinite population with iid data and the problem we study here is that when, as is common, the probability sample is obtained using sampling without replacement, Sample A data are not iid.
@@
We build on the work of Chen, Li and Wu (2020)[6] (henceforth, CLW). CLW developed a DR estimator for $\bar{Y}$ using parametric models for the nuisance functions. Yang et al. (2020)[29] extended this method to allow for variable selection by smoothly clipped absolute deviation (SCAD). We generalise the work of CLW (and Yang et al.) to allow for general data-adaptive estimation of the nuisance functions. Our approach uses cross-fitting, a technique popular in the field of debiased machine learning.
@@
The structure of the article is as follows. In Section 2, we describe our design-model-based inference framework. In Section 3, we describe a von Mises expansion and cross-fitting, both fundamental to our approach. In Section 4, we consider the scenario where clusters (or ‘primary sampling units’) are sampled using simple random sampling (i.e. with equal probabilities) without replacement and then individuals within clusters are sampled using an arbitrary (possibly multi-stage) sampling design that obeys the independence and invariance conditions[24]. Section 5 considers the more general scenario where the cluster sampling probabilities vary. In Section 6, we describe a ratio-style DR estimator of $\bar{Y}$ that is typically more efficient than the DR estimator described in earlier sections. Section 7 covers targeted maximum likelihood estimators (TMLEs), which are an alternative to the estimating equations-style DR estimators described in earlier sections. Possible data-adaptive estimators of the nuisance functions are discussed in Section 8. In Section 9, we present a simulation study that compares DR methods using data-adaptive estimators of nuisance functions with those that use parametric estimators. In this article, we do not use data on the $Y$ values of individuals in Sample A. When such data are available, $\bar{Y}$ can be estimated using only Sample A, without needing Sample B. However, efficiency may be gained by combining the DR estimator that uses the $Y$ values only of individuals in Sample B with the estimator that uses only Sample A[18, 13, 25]. This, along with other issues, is discussed in Section 10.
2 Data-generating mechanism and estimator
@@
The model for drawing Sample A from this finite population is as follows. The variable $R^{A}_{ji}$ will be a binary variable indicating whether individual $i$ in cluster $j$ ($i=1,\ldots,N_{j};\;j=1,\ldots,J$) is included in Sample A. First, a sample of size $M$ ($M<J$) clusters is drawn from the $J$ clusters in the population, with the probability that cluster $j$ is included in the sample being proportional to $h(N_{j},X_{j1},\ldots,X_{jN_{j}})$ for some function $h$ of $N_{j}$ and $X_{j1},\ldots,X_{jN_{i}}$. A possible example would be $h(N_{j},X_{j1},\ldots,X_{jN_{j}})=N_{j}$. Let $R^{C}_{j}=1$ if cluster $j$ is drawn, and $R^{C}_{j}=0$ otherwise. If $R^{C}_{j}=0$, then $R^{A}_{ji}=0$ ($i=1,\ldots,N_{j}$). If $R^{C}_{j}=1$, then $(R^{A}_{j1},\ldots,R^{A}_{jN_{j}})$ is drawn according to some (possibly multistage) sampling design that can depend on $N_{j}$ and $(X_{j1},\ldots,X_{jN_{j}})$ but which, given $N_{j}$, $(X_{j1},\ldots,X_{jN_{j}})$ and $R^{C}_{j}=1$, does not depend on $(Y_{j1},\ldots,Y_{jN_{j}})$, $\{(N_{k},X_{k1},\ldots,X_{kN_{k}},Y_{k1},\ldots,Y_{kN_{k}}):\;k=1,\ldots,j-1,j+1,\ldots J\}$ or $(R^{C}_{1},\ldots,R^{C}_{j-1},R^{C}_{j+1},\ldots,R^{C}_{J}$). This sampling of $(R^{A}_{j1},\ldots,R^{A}_{jN_{j}})$ is done independently of the sampling of $(R^{A}_{k1},\ldots,R^{A}_{kN_{k}})$ within any other cluster $k$ with $R^{C}_{k}=1$. Note this means we are assuming independence and invariance (see pp134–135 of [24]) for the sampling design. Let $\pi^{C}_{j}=P(R^{C}_{j}=1\mid\mathcal{F}_{J})$ and $\pi^{A\mid C}_{ji}=P(R^{A}_{ji}=1\mid R^{C}_{j}=1,N_{j},X_{j1},\ldots,X_{jN_{j}})$. Let $\pi^{A}_{ji}=\pi^{C}_{j}\times\pi^{A\mid C}_{ji}$ denote the first-order inclusion probability for individual $i$ in cluster $j$. For simplicity, we do not consider stratified sampling here, but this is discussed in Section 10.
@@
Now relabel the $\sum_{j=1}^{J}N_{j}$ individuals so that, for each $j=1,\ldots,J$, the $N_{j}$ individuals in cluster $j$ are labelled as individuals $\sum_{k=1}^{j-1}N_{j}+1,\sum_{k=1}^{j-1}N_{j}+2,\ldots,\sum_{k=1}^{j}N_{j}$. Let $D_{i}$ ($1\leq D_{i}\leq J$) denote the number of the original cluster to which individual $i$ belongs. Finally, let $n=n(\mathcal{F}_{J})=\sum_{j=1}^{J}N_{j}$ denote the total number of individuals in the population. We can now write the population mean as $\bar{Y}=\frac{1}{n}\sum_{i=1}^{n}Y_{i}$.
The observed data are
$\{(R^{A}_{i},R^{A}_{i}X_{i},R^{A}_{i}\pi^{A}_{i},R^{A}_{i}D_{i},R^{B}_{i},R^{B}_{i}X_{i},R^{B}Y_{i},R^{B}_{i}D_{i});\;i=1,\ldots,n\},$
i.e. the $X$ values and sampling probabilities of individuals in Sample A, the $X$ and $Y$ values of individuals in Sample B, and the index of the cluster to which each of these individuals belongs. Note that our notation suggests that we know whether an individual in Sample A also appears in Sample B, and vice versa. In fact, this is unnecessary; this notation is used only for convenience. We shall, however, need to know, for each individual in Sample A and each individual in Sample B, whether they belong to the same cluster (this is used in the cross-fitting procedure — see Section 3).

### v2

@@ 1 Introduction
Several methods have recently been proposed for using data from a probability sample (Sample A) as auxiliary information to address nonrepresentative of a nonprobability sample (Sample B). These methods fall into three classes, all of which require data on some covariates $X$ that are observed on the individuals in both samples. The first class of methods, which include inverse probability weighting (IPW) methods, involve estimating weights for the individuals in Sample B such that the weighted Sample B is representative of the population. The second class, known as mass imputation methods, involve using the relation between $X$ and $Y$ in Sample B to impute the $Y$ values in Sample A. The third class, known as double robust (DR) methods, combine IPW and mass imputation. IPW methods require a consistent estimator of the conditional probability that an individual belongs to Sample B given $X$. Mass imputation methods require a consistent estimator of the conditional expectation of $Y$ given $X$. We shall refer to this conditional probability and expectation as ‘nuisance’ functions, because they are not of direct interest. DR methods use estimators of both nuisance functions but only require one of these estimators to be consistent.
@@
A parametric model could be used for each of the two nuisance functions, but the resulting estimator would typically be inconsistent if that parametric model — or, in the case of DR estimators, both parametric models — were misspecified. There is therefore considerable interest in using more flexible ‘data-adaptive’ or ‘machine-learning’ estimators of the nuisance functions in order to minimise this risk of model misspecification. A number of researchers have done this. Ferri-Garcia et al. (2020, 2024)[17, 18], Rueda et al. (2023)[14] and Rueda et al. (2024)[13] use IPW and estimate the conditional probability nuisance function with k-nearest neighbours, random forest or generalised boosting models (specifically, XGBoost). Castro-Martin et al. (2020)[2] use IPW or mass imputation, and estimate the conditional probability or expectation nuisance function using random forest or XGBoost. Ferri-Garcia et al. (2022)[16] use IPW and estimate the nuisance function using classification and regression trees (CART). Castro-Martin et al. (2021)[3] use IPW, mass imputation or a DR estimator, and estimate the nuisance functions (or functions) using XGBoost. Cobo et al. (2025)[12] use a DR estimator and estimate the nuisance functions using XGBoost, k-nearest neighbours or neural nets. To reduce the variability of the estimated inverse probability weights, Castro-Martin et al. (2021)[3], Rueda et al. (2024)[13] and Cobo et al. (2025)[12] apply kernel smoothing to them. Wang et al. (2020,2022)[28, 29] developed a theoretical justification for this kernel smoothing approach when logistic regression is used, but not when data-adaptive methods are used. Chen et al. (2022)[5] use mass imputation with kernel smoothing, show that the resulting estimator of $\bar{Y}$ is asymptotically normal and provide a formula for the variance. They note, however, that this method is subject to the curse of dimensionality. Chlebicki et al. (2024)[10] use mass imputation with predictive mean matching. This matching provides some limited robustness to misspecification of a parametric model for the nuisance conditional expectation function, because it requires only that the true conditional expectation be a monotone function of the limiting value of the estimated conditional expectation. Chlebicki et al. suggest that a data-adaptive method could be used to estimate this nuisance function, but do not prove that the resulting estimator of $\bar{Y}$ would be $n^{1/2}$-consistent.
@@
A closely related problem to that of estimating $\bar{Y}$ for a finite population by combining data from probability and nonprobability samples is that of estimating $E(Y)$, the expectation of $Y$ in an infinite population, when data $(X_{i},Y_{i})$ $(i=1,\ldots,n)$ are independent and identically distributed (iid), $X$ is fully observed and $Y$ is missing at random given $X$. IPW, regression imputation and DR methods exist for this problem. The first two require consistent estimators of, respectively, the conditional probability that $Y$ is observed given $X$ and the conditional expectation of $Y$ given $X$. DR methods involve estimating both these nuisance functions, but require only that one of them is estimated consistently. For IPW and imputation, it is known that using data-adaptive methods to estimate the nuisance function can cause problems. Data-adaptive methods typically yield nuisance function estimators that converge slowly, and this slow convergence is inherited by the IPW or imputation estimator of $E(Y)$, which can then have considerable finite-sample bias; confidence intervals can also have poor coverage[27]. The DR approach (also known as ‘debiased machine learning’) can overcome this issue, because DR estimators do not inherit the slow convergence of the nuisance function estimators[27, 15, 24, 21, 8, 7, 9]. Given the similarity of the problem of estimating $E(Y)$ for an infinite population with iid data to the problem of estimating $\bar{Y}$ for a finite population using a nonprobability sample, there is reason to be concerned that the same issue may well also apply to the latter problem. In this article, we study the use of a DR estimator of $\bar{Y}$ for a finite population when combining data from a probability sample and a nonprobability sample and using data-adaptive estimators of the nuisance functions. Our goal is to provide a theoretical justification for the validity of this approach. An important difference between the problem of estimating $E(Y)$ for an infinite population with iid data and the problem we study here is that when, as is common, the probability sample is obtained using sampling without replacement (SWOR), Sample A data are not independent.
@@
We build on the work of Chen, Li and Wu (2020)[6] (henceforth, CLW). CLW developed a DR estimator for $\bar{Y}$ using parametric models for the nuisance functions. Yang et al. (2020)[30] extended this method to allow for variable selection by smoothly clipped absolute deviation. We generalise the work of CLW (and Yang et al.) to allow for general data-adaptive estimation of the nuisance functions. Our approach uses cross-fitting, a technique popular in the field of debiased machine learning.
@@
The structure of the article is as follows. In Section 2, we describe our assumed data-generating mechanism and design-model-based inference framework. In Section 3, we describe a DR estimator for $\bar{Y}$, its von Mises expansion and cross-fitting, which is fundamental to our approach. In Section 4, we consider the scenario where clusters (or ‘primary sampling units’) are sampled using simple random sampling without replacement (SRSWOR) and then individuals within clusters are sampled using an arbitrary (possibly multi-stage) sampling design that obeys the independence and invariance conditions[25]. Section 5 considers the more general scenario where the cluster sampling probabilities vary. In Section 6, we describe a ratio-type DR estimator of $\bar{Y}$ that is typically more efficient than the DR estimator described in earlier sections. Section 7 covers targeted maximum likelihood estimators (TMLEs), which are an alternative to the estimating equations-type DR estimators described in earlier sections. Possible data-adaptive estimators of the nuisance functions are discussed in Section 8. In Section 9, we present a simulation study that compares DR methods using data-adaptive estimators of nuisance functions with those that use parametric estimators. In this article, we do not use data on the $Y$ values of individuals in Sample A. When such data are available, $\bar{Y}$ can be estimated using only Sample A, without needing Sample B. However, efficiency may be gained by combining the DR estimator that uses the $Y$ values only of individuals in Sample B with the estimator that uses only Sample A[19, 14, 26]. This, along with other issues, is discussed in Section 10.
2 Assumed data-generating mechanism
@@
The model for drawing Sample A from this finite population is as follows. The variable $R^{A}_{ji}$ will be a binary variable indicating whether individual $i$ in cluster $j$ ($i=1,\ldots,N_{j};\;j=1,\ldots,J$) is included in Sample A. First, a sample of size $M$ ($M<J$) clusters is drawn without replacement from the $J$ clusters in the population, with the probability that cluster $j$ is included in the sample being proportional to $h(N_{j},X_{j1},\ldots,X_{jN_{j}})$ for some function $h$ of $N_{j}$ and $X_{j1},\ldots,X_{jN_{i}}$. An example would be $h(N_{j},X_{j1},\ldots,X_{jN_{j}})=N_{j}$. Let $R^{C}_{j}=1$ if cluster $j$ is drawn; $R^{C}_{j}=0$ otherwise. If $R^{C}_{j}=0$, then $R^{A}_{ji}=0$ ($i=1,\ldots,N_{j}$). If $R^{C}_{j}=1$, then $(R^{A}_{j1},\ldots,R^{A}_{jN_{j}})$ is drawn according to some (possibly multistage) SWOR design that can depend on $N_{j}$ and $(X_{j1},\ldots,X_{jN_{j}})$ but which, given $N_{j}$, $(X_{j1},\ldots,X_{jN_{j}})$ and $R^{C}_{j}=1$, does not depend on $(Y_{j1},\ldots,Y_{jN_{j}})$, $\{(N_{k},X_{k1},\ldots,X_{kN_{k}},Y_{k1},\ldots,Y_{kN_{k}}):\;k=1,\ldots,j-1,j+1,\ldots J\}$ or $(R^{C}_{1},\ldots,R^{C}_{j-1},R^{C}_{j+1},\ldots,R^{C}_{J}$). This sampling of $(R^{A}_{j1},\ldots,R^{A}_{jN_{j}})$ is done independently of the sampling of $(R^{A}_{k1},\ldots,R^{A}_{kN_{k}})$ within any other cluster $k$ with $R^{C}_{k}=1$. Note this means we are assuming independence and invariance (see pp134–135 of [25]) for the sampling design. Let $\pi^{C}_{j}=P(R^{C}_{j}=1\mid\mathcal{F}_{J})$ and $\pi^{A\mid C}_{ji}=P(R^{A}_{ji}=1\mid R^{C}_{j}=1,N_{j},X_{j1},\ldots,X_{jN_{j}})$. Let $\pi^{A}_{ji}=\pi^{C}_{j}\times\pi^{A\mid C}_{ji}$ denote the first-order inclusion probability for individual $i$ in cluster $j$. For simplicity, we do not consider stratified sampling here, but this is discussed in Section 10.
@@
Now relabel the $\sum_{j=1}^{J}N_{j}$ individuals so that, for each $j=1,\ldots,J$, the $N_{j}$ individuals in cluster $j$ are labelled as individuals $\sum_{k=1}^{j-1}N_{k}+1,\sum_{k=1}^{j-1}N_{k}+2,\ldots,\sum_{k=1}^{j}N_{k}$. Let $D_{i}$ ($1\leq D_{i}\leq J$) denote the number of the cluster to which individual $i$ belongs. Finally, let $n=n(\mathcal{F}_{J})=\sum_{j=1}^{J}N_{j}$ denote the total number of individuals in the population. We can now write the population mean as $\bar{Y}=n^{-1}\sum_{i=1}^{n}Y_{i}$.
The observed data are the $X$ values and sampling probabilities $\pi^{A}$ of individuals in Sample A, the $X$ and $Y$ values of individuals in Sample B, and the index of the cluster to which each of these individuals belongs, i.e.
$\{(R^{A}_{i},R^{A}_{i}X_{i},R^{A}_{i}\pi^{A}_{i},R^{A}_{i}D_{i},R^{B}_{i},R^{B}_{i}X_{i},R^{B}Y_{i},R^{B}_{i}D_{i}):\;i=1,\ldots,n\}.$
Note that our notation suggests that we know whether an individual in Sample A also appears in Sample B, and vice versa. In fact, this is unnecessary; this notation is used only for convenience. We shall, however, need to know, for each individual in Sample A and each individual in Sample B, whether they belong to the same cluster (this information is used in the cross-fitting procedure — see Section 3).
We shall consider the asymptotic repeated-sampling distributions of several DR estimators of $\bar{Y}$ conditional on $\mathcal{F}_{J}$ in an asymptotic framework in which $J\rightarrow\infty$ and $M\rightarrow\infty$ with $M/J$ converging to a constant. This means we do not need to distinguish between $M\rightarrow\infty$ and $J\rightarrow\infty$. The data-generating models for drawing Samples A and B from the finite population do not change as $M\rightarrow\infty$.
3 DR estimator and its von Mises expansion
3.1 iid data and infinite population
It is instructive to consider the simpler context of estimating $E(Y)$ for an infinite population using a sample of $n$ iid individuals with $X$ observed for all $n$ individuals, $Y$ observed for only some of them, and these data being missing at random. In this context, a DR estimator for $E(Y)$ is $\hat{\theta}_{\rm iid}(\hat{\pi}_{\rm iid},\hat{m}_{\rm iid})$, where
$\hat{\theta}_{\rm iid}(\pi_{\rm iid},m_{\rm iid})=\frac{1}{n}\sum_{i=1}^{n}m_{\rm iid}(X_{i})+\frac{R_{i}}{\pi_{\rm iid}(X_{i})}\{Y_{i}-m_{\rm iid}(X_{i})\},$
$R$ is an indicator that $Y$ is observed, $\hat{\pi}_{\rm iid}(X)$ is an estimator of $\pi_{\rm iid0}(X)=P(R=1\mid X)$ and $\hat{m}_{\rm iid}(X)$ is an estimator of $m_{\rm iid0}(X)=E(Y\mid X)$. Hines et al. (2022)[21] describe how a von Mises expansion can be used to show that if $\hat{\pi}_{\rm iid}$ and $\hat{m}_{\rm iid}$ converge to $\pi_{\rm iid0}$ and $m_{\rm iid0}$ sufficiently fast as $n\rightarrow\infty$, then $\hat{\theta}_{\rm iid}(\hat{\pi}_{\rm iid},\hat{m}_{\rm iid})=\hat{\theta}_{\rm iid}(\pi_{\rm iid0},m_{\rm iid0})+o_{p}(n^{-1/2})$, and so $\hat{\theta}_{\rm iid}(\hat{\pi}_{\rm iid},\hat{m}_{\rm iid})$ has the same asymptotic normal distribution as $\hat{\theta}_{\rm iid}(\pi_{\rm iid0},m_{\rm iid0})$. The required convergence rates of $\hat{\pi}_{\rm iid}$ and $\hat{m}_{\rm iid}$ are slower than the parametric $n^{1/2}$ rate, which enables the use of data-adaptive estimators, provided that either these estimators satisfy the Donsker condition or cross-fitting is used.
3.2 Sampling without replacement from a finite population
In this article, we obtain an analogous result for our sample survey setting. We make use of cross-fitting. Randomly partition the set of $n$ individuals into $K$ (e.g. $K=5$) subsets, called folds. The way this is done will be described in Sections 4 and 5. Let $\mathcal{S}_{k}\subset\{1,\ldots,n\}$ denote the set of indices of the individuals belonging to fold $k$ ($k=1,\ldots,K$), and let $n_{k}=|\mathcal{S}_{k}|$ denote the number of these individuals. Let $\mathcal{S}_{-k}$ denote the indices of the individuals not in fold $k$ (so, $\mathcal{S}_{k}\cup\mathcal{S}_{-k}=\{1,\ldots,n\}$). Let $\hat{m}_{k}$ be an estimator of $m_{0}$ calculated using only the data $\{(X_{i},Y_{i}):R^{B}_{i}=1\mbox{ and }i\in\mathcal{S}_{-k}\}$. Let $\hat{\pi}^{B}_{k}$ be an estimator of $\pi^{B}_{0}$ obtained using only the data $\{(R^{A}_{i},R^{B}_{i},R^{A}_{i}\pi^{A}_{i},R^{A}_{i}X_{i},R^{B}_{i}X_{i}):i\in\mathcal{S}_{-k}\}$. We discuss possible estimators in Section 8. Write $\underline{\pi}^{B}=(\pi^{B}_{1},\ldots,\pi^{B}_{K})$, $\underline{m}=(m_{1},\ldots,m_{K})$, $\underline{\hat{\pi}}^{B}=(\hat{\pi}^{B}_{1},\ldots,\hat{\pi}^{B}_{K})$ and $\underline{\hat{m}}=(\hat{m}_{1},\ldots,\hat{m}_{K})$.